# Data for Mathematical Copilots: Better Ways of Presenting Proofs for Machine Learning

### v1

@@ Abstract
The suite of datasets commonly used to train and evaluate the mathematical capabilities of AI-based mathematical copilots (primarily large language models) exhibit several shortcomings. These limitations include a restricted scope of mathematical complexity, typically not exceeding lower undergraduate-level mathematics, binary rating protocols and other issues, which makes comprehensive proof-based evaluation suites difficult. We systematically explore these limitations and contend that enhancing the capabilities of large language models, or any forthcoming advancements in AI-based mathematical assistants (copilots or “thought partners”), necessitates a paradigm shift in the design of mathematical datasets and the evaluation criteria of mathematical ability: It is necessary to move away from result-based datasets (theorem statement to theorem proof) and convert the rich facets of mathematical research practice to data LLMs can train on. Examples of these are mathematical workflows (sequences of atomic, potentially subfield-dependent tasks that are often performed when creating new mathematics), which are an important part of the proof-discovery process. Additionally, we advocate for mathematical dataset developers to consider the concept of “motivated proof”, introduced by G. Pólya in 1949, which can serve as a blueprint for datasets that offer a better proof learning signal, alleviating some of the mentioned limitations. Lastly, we introduce math datasheets for datasets, extending the general, dataset-agnostic variants of datasheets: We provide a questionnaire designed specifically for math datasets that we urge dataset creators to include with their datasets. This will make creators aware of potential limitations of their datasets while at the same time making it easy for readers to assess it from the point of view of training and evaluating mathematical copilots.
@@ 1 AI Systems for Mathematicians – Present and Future
There has been a rapid surge in research around AI for mathematics. For example, AlphaGeometry was shown to solve mathematical geometry problems on the level of the International Math Olympiad (IMO) (Trinh et al., 2024), Numina won the first AIMO Progress Prize, [^1] https://aimoprize.com/updates/2024-07-20-progress-prize-results by providing a model capable of turning mathematical questions into executable Python code, and previously GPT-4 (OpenAI, 2023) was shown to attain, in some cases, the performance of an undergraduate university student (Frieder et al., 2023a). High accuracies on standard mathematics benchmarks (Reid et al., 2024) suggest that many benchmarks may be close to being solved.
@@
    some datasets being overstudied (e.g., the many versions of the GSM8K dataset (Cobbe et al., 2021)), whereas other data-related aspects (few datasets comprising advanced mathematics, or reflecting tool use in mathematics) are neglected;
@@
    Highly specialized, “narrow” models that fulfill all of the criteria of 1) being designed for a precisely defined mathematical domain or mathematical problem, 2) requiring domain knowledge to be used (such as formal languages as input), 3) that cannot explain themselves, and 4) do not allow free-form interaction. Examples of such specialized models vary widely in terms of how their foundations look like. Specialized models could rely on symbolic reasoners, such as the recent solver for elementary geometry (Newclid (Sicca et al., 2024) which supersedes AlphaGeometry (Trinh et al., 2024)) or for inequality problems ((Wei et al., 2024)). Olympiad-level mathematics problems are within reach of both. Alternatively, such specialized models could simply be regression models, relying on raw mathematical data to infer new relationships between mathematical objects, such as for knot theory or representation theory (Davies et al., 2021), or group theory (He et al., 2023). Many further approaches exist (Romera-Paredes et al., 2024; Fawzi et al., 2022). These systems typically require a significant amount of specialization by a user, such as mastering a formal input language (e.g., AlphaGeometry) or understanding the machine learning model itself (e.g., (Davies et al., 2021)) beyond the necessary mathematical domain knowledge.
@@
    General purpose, “broad” models that can interact with humans via natural language, understand input images, and use tools (being able to call, in particular, specialized models, which can be viewed as a particular kind of “tool”, as we argue below), can provide general assistance to a mathematician (both with routine tasks, such as literature search, and domain-dependent tasks, such as counterexample search) and are not tied to a specific area of mathematics. These are what we refer to as “Mathematical Copilots”, as their goal is not simply to provide correct proofs but also to enhance the user’s understanding of the results and guide them towards their own discoveries (Collins et al., 2024a). Currently, the best candidates for these systems are large language models (LLMs), which we take to include multi-modal models that also admit images as components of the prompt. We contend that LLMs represent early predecessors of future general-purpose models (mathematical copilots), ultimately acting as mathematicians’ “thought partners” (Collins et al., 2024b). Anecdotal evidence of such usefulness of current LLMs has been presented in various discussion forums, [^2] https://www.reddit.com/r/math/comments/14p6j5c/terence_tao_on_using_gpt4_to_help_with_math/ while noting the failures that LLMs often succumb to. However, there are concrete ways in which LLMs can be improved to become mathematical copilots: As indicated by the tool-integrated-reasoning approach, championed by models such as ToRA (Gou et al., 2024) or Numina, [^3] https://huggingface.co/AI-MO/NuminaMath-7B-TIR general LLMs can incorporate other symbolic systems to which they can delegate symbolic tasks to. The mixture-of-experts paradigm (Cai et al., 2024), on the other hand, highlights different ways in which LLMs can incorporate other LLMs as specially trained “modules” to which they can delegate mathematical tasks that are not symbolic in nature but still require particular mathematical skills. Such a hierarchical system can help mathematicians both on routine academic tasks (see (Frieder et al., 2023b) for a short, high-level overview) and specialized ones.
    We envision that such models will require significantly less effort from users to use them, compared to specialized models, being self-explanatory, where natural language is the main mode of interaction. (While the outlined architectural approaches provide a viable step forward, in some cases, no data exists on which to train these models; see below.)
@@
    Fully automated, “universal” models that can generate mathematical theorems and theories autonomously, merely by being directed to a mathematical result. We envision that these systems will be the true successors of automated theorem provers (ATPs) such as the Vampire (Riazanov and Voronkov, 2002; Kovács and Voronkov, 2013) or E (Schulz, 2002) ATPs. These systems have not been widely adopted by the mathematical community. While there have been attempts to imbue these systems with (non-LLM) machine learning techniques (Holden, 2021), these have not yet dramatically increased their performance. Nonetheless, once general-purpose models have been established, it is conceivable one could put these in automation loops to obtain “AI mathematicians”. For the domain of AI, this has already been recently investigated in the form of an “AI Scientist” (Lu et al., 2024a) that autonomously generated machine learning research articles. These are systems that conceivably will require minimal human input, merely guiding them towards certain theorems that one wants to see analyzed (proved, refuted, with comments on whether slightly different variations of a theorem, with slightly altered hypotheses, admit a proof).
We focus in this article on general purpose models, as outlined above, which are mathematical copilots. LLMs are currently the systems that hold the most promise to become, given a better data foundation, general purpose models. Fully automated, universal models are currently out of reach, as they are contingent on the former category of models. These are the systems that are ultimately sought when one speaks of “AI Mathematicians” (Bengio and Malkin, 2024). However, without the proper datasets, it is unlikely that these will come into existence. Hence, our focus is on the data that is needed to arrive at general-purpose models, the mathematical copilots. We have deliberately avoided drawing a hard boundary among the properties a model must fulfill to be placed in one of these three because the boundaries will be fluid, as the levels of automation and required human interaction are continuous scales.
@@
Our main reader audience consists of machine learning researchers, as we intend this article to raise awareness of these issues. Automated theorem provers (ATPs) and interactive theorem provers (ITPs), which have a decades-old history (Harrison et al., 2014), also had the ambition of realizing the goal of implementing mathematical copilots. Unfortunately, history has shown that if the computer science community and the mathematical community do not operate in lockstep, there is a risk of obtaining systems that will not be of interest to practicing research mathematicians; maybe controversially, Blanchette et al. (2012) mentions that certain efforts to make these systems more user-friendly have perhaps reached a “plateau”. (Nonetheless, ATPs and ITPs have, in other domains, contributed significantly to driving the field of software automation forward.)
@@ 2 An Overview of Data-Related Issues
Without using tools specialized for mathematics, the current general models, LLMs, simply by well-designed training methodologies and trained on mathematical data, have demonstrated unprecedented capabilities in generating humanlike mathematical text, solving complex problems, and even engaging in creative problem-solving. For example, GPT-4 has performed promisingly on undergraduate-level mathematics formulated in natural language, as pointed out by Frieder et al. (2023a). Math-Specialized Gemini 1.5 Pro (Reid et al., 2024), a commercial model by Google not available to the public, has been reported to have an accuracy of over 90% on the MATH dataset (Hendrycks et al., 2021), one of the most widely-used datasets for testing mathematical ability. This score has recently been replicated by an open-weight model, QwQ. [^4] https://qwenlm.github.io/blog/qwq-32b-preview/ Attaining such a high score is equivalent to achieving the ability of an IMO gold medallist (according to (Hendrycks et al., 2021)). However, to date, no reports have been made where the mathematical benchmarks of QwQ are tested for contamination; not undertaking tests may lead to this score not being reproducible on a different dataset from the same difficulty and problem type distribution (Yang et al., 2023; Xu et al., 2024). Further notable general-purpose LLMs are open-weight models with strong reported performance on baseline datasets, such as MATH and GSM8K (Cobbe et al., 2021), the DeepSeek family of models (Shao et al., 2024; Liu et al., 2024; Zhu et al., 2024; Xin et al., 2024), and the Qwen family of models (Bai et al., 2023; Yang et al., 2024a; Hui et al., 2024; Yang et al., 2024b). Other LLMs focus more on specific abilities such as MathPrompter (Imani et al., 2023), which associates a confidence value to arithmetic problem solutions, MathVista (Lu et al., 2024b), which solves geometric reasoning problems, WizardMath (Luo et al., 2023), or Llemma (Azerbayev et al., 2024), which was specialized through further pretraining on a math corpus, and is both open-weight and open-source. A special use case where LLMs have found success is in interactive theorem provers (ITPs) such as Isabelle (Nipkow et al., 2002) or Lean (de Moura and Ullrich, 2021), where they are used to prove theorems most prominently in the context of generating proofs of given formal statements (First et al., 2023; Zheng et al., 2023; Wang et al., 2024a), autoformalization (Szegedy, 2020; Jiang et al., 2023), or providing code snippets of formal mathematics (Song et al., 2024a). For a survey of deep learning approaches on both formal and informal (natural language) mathematics, we refer to Lu et al. (2023), and for more recent information on language models, including their performance on mainstream datasets, we refer to Zhao et al. (2023).
@@
This paper argues that the current dataset landscape does not support the advancement of such general-purpose AI systems, in particular LLMs, towards a level of mathematical performance that makes them usable as daily mathematical thought partners (Collins et al., 2024b) that help mathematicians push the boundaries of what is known, and capture the richness of mathematical reasoning and invention that human mathematical minds are capable of (Zhang et al., 2023a; Dehaene, 2011; Feigenson et al., 2004). Our focus thus includes, in particular, research-level mathematics, which involves many steps beyond deriving or formalizing proofs – the aspects of mathematical practice that are currently most strongly represented in data.
@@
The spectrum of tasks a copilot can assist with is not solely related to mathematics: programmers presently carry out programming tasks, such as fixing bugs or other GitHub issues (Jimenez et al., 2024), using various code-generating copilots, such as the GitHub Copilot (Chen et al., 2021); the emerging Lean Copilot ecosystem (Yang et al., 2024c; Song et al., 2024b) assists users in the task of formalizing a mathematical theorem; Tutor CoPilot (Wang et al., 2024b) provides real-time guidance to human tutors that teach school students mathematics at the level of math word problems. Yet, mathematics, with its rich set of domain-dependent workflows, tools, and modes of thinking, requires copilots with exceptionally broad capabilities. A mathematical copilot that is useful to a mathematician must significantly exceed the capabilities of the three exemplary copilots mentioned above, which are systems that are designed to help with a narrow technical set of tasks.
@@
To give an overview of our contributions in that section, current natural language datasets typically model a question-response interaction, such as posing a problem and writing a solution. We identify four dimensions where such datasets frequently fall short of being able to assess advanced mathematics:
@@
  • Standardized Interaction Modes: Often, natural language datasets only contain question-answer mathematical problems. The daily experience of a mathematician is much richer than that (Zhang et al., 2023a), since to prove/search for a theorem, all kinds of “lateral” modes of thinking have to be applied. Examples are intermediate conjecture formulation, (counter)example search, and finding analogies with other theorems. These are typically not found in textbooks, so there is no representation of these in terms of data. We point to  Collins et al. for an exposition of the problems of static evaluation for natural-language mathematics problems in practice.

### v2

@@ Abstract
The datasets and benchmarks commonly used to train and evaluate the mathematical capabilities of AI-based mathematical copilots (primarily large language models) exhibit several shortcomings and misdirections. These range from a restricted scope of mathematical complexity to limited fidelity in capturing aspects beyond the final, written proof (e.g. motivating the proof, or representing the thought processes leading to a proof). These issues are compounded by a dynamic reminiscent of Goodhart’s law: as benchmark performance becomes the primary target for model development, the benchmarks themselves become less reliable indicators of genuine mathematical capability. We systematically explore these limitations and contend that enhancing the capabilities of large language models, or any forthcoming advancements in AI-based mathematical assistants (copilots or “thought partners”), necessitates a course correction both in the design of mathematical datasets and the evaluation criteria of the models’ mathematical ability. In particular, it is necessary for benchmarks to move beyond the existing result-based datasets that map theorem statements directly to proofs, and instead focus on datasets that translate the richer facets of mathematical research practice into data that LLMs can learn from. This includes benchmarks that supervise the proving process and the proof discovery process itself, and we advocate for mathematical dataset developers to consider the concept of “motivated proof”, introduced by G. Pólya in 1949, which can serve as a blueprint for datasets that offer a better proof learning signal, alleviating some of the mentioned limitations.
@@ 1 AI Systems for Mathematicians – Present and Future
There has been a rapid surge in research around AI for mathematics. For example, AlphaGeometry was shown to solve mathematical geometry problems on the level of the International Math Olympiad (IMO) (trinh2024solving), Numina won the first AIMO Progress Prize, [^1] https://aimoprize.com/updates/2024-07-20-progress-prize-results by providing a model capable of turning mathematical questions into executable Python code, and previously GPT-4 (openai2023gpt4) was shown to attain, in some cases, the performance of an undergraduate university student (frieder2023mathematical). High accuracies on standard mathematics benchmarks (reid2024gemini) suggest that many benchmarks may be close to being solved.
Update since the December 2024 version.
This manuscript was initially circulated in December 2024. Since then, reported performance on boxed-answer benchmarks has continued to increase, and some reasoning-focused systems report near-saturation on MATH-style evaluations (deepseek2025r1; qwen2024qwq32bpreview). At the same time, many of the issues discussed in this paper have seen fewer changes than model scores: public math benchmarks remain largely static, scoring often still collapses rich outputs to a single correctness bit, and contamination risk persists as benchmarks diffuse into training corpora. The revisions in this version therefore focus on incorporating recent benchmark results while keeping the central claim unchanged: progress toward mathematical copilots depends on data and evaluation that represent intermediate reasoning and workflow competence, not only final answers.
@@
    some datasets being overstudied (e.g., the many versions of the GSM8K dataset (cobbe2021training)), whereas other data-related aspects (few datasets comprising advanced mathematics or reflecting tool use in mathematics) are neglected;
@@
These problems are not unique to AI for mathematics, but the broader machine learning community has begun systematic self-assessment efforts that the mathematical AI community has yet to undertake. For instance, bhardwaj2024curation present an evaluation framework grounded in data curation principles from librarianship and archives, and use it to assess 60 datasets from the NeurIPS Datasets and Benchmarks track. Their analysis shows gaps in documentation around ethical considerations, data management, and environmental footprint, and they provide concrete recommendations for improving peer-review standards. No comparable audit exists for mathematical datasets. Such an effort would be valuable: mathematical datasets come with domain-specific concerns (e.g., difficulty calibration, formal-informal alignment, contamination risks) that generic ML dataset assessments do not address.
@@
    Highly specialized, “narrow” models that fulfill all of the criteria of 1) being designed for a precisely defined mathematical domain or mathematical problem, 2) requiring domain knowledge to be used (such as formal languages as input), 3) that cannot explain themselves, and 4) do not allow free-form interaction. Examples of such specialized models vary widely in terms of how their foundation is set up (model architecture, training data), and in what they assist mathematicians. Specialized models could rely on symbolic reasoners, such as the recent solver for elementary geometry (Newclid (sicca2024newclid), which supersedes AlphaGeometry (trinh2024solving)) or for inequality problems ((wei2024proving)). Olympiad-level mathematics problems are within reach of both, and these assist mathematicians solely in solving closed-ended questions. Alternatively, such specialized models could simply be regression models, relying on raw mathematical data to infer new relationships between mathematical objects, such as for knot theory or representation theory (davies2021advancing) or group theory (he2023learning), assisting mathematicians in finding conjectures (which then need to be proved by hand). Many further approaches exist (romera2024mathematical; fawzi2022discovering; hashemi2024can), see Table 1 for more information. These systems typically require a significant amount of specialization by a user, such as mastering a formal input language (e.g., AlphaGeometry and Newclid, which both require mastery of the JGEX formal language) or understanding the machine learning model itself (e.g., (davies2021advancing)) beyond the necessary mathematical domain knowledge. The data on which they operate also varies strongly: Symbolic solvers rely on a symbolic language; models like (davies2021advancing) rely on tabular data, while models like (hashemi2024can) combine different data types.
    System Reference Architecture class How it helps mathematicians User understanding
    Closed-ended problem solving (verifiable / near-verifiable)
    AlphaGeometry trinh2024solving Hybrid (neural + symbolic) Solves Euclidean geometry problems via learned components plus symbolic deduction; uses a domain-specific formal encoding (JGEX). IF
    Newclid sicca2024newclid Symbolic (search / ATP-style) Solves Euclidean geometry problems in JGEX, producing checkable derivations. IF
    Inequality prover wei2024proving Symbolic (automation) Automates structured inequality proving in restricted families, producing derivations that can often be mechanically verified. IF
    Conjecture, counterexample, and construction discovery (requires downstream validation)
    Graph-building RL counterexamples Wagner:2021Constructing RL policy (often MLP/GNN) Finds counterexamples by constructing discrete objects in an environment with an automatically scored objective. PL
    PatternBoost Charton:2024Patternboost Transformer + local search Discovers combinatorial constructions by alternating neural generation with search, targeting automatically checkable properties. PL
    Invariant-correlation models (e.g., knot theory) davies2021advancing Supervised regression (often linear / tree-based) Suggests conjectural relationships between computable invariants; outputs hypotheses that mathematicians then prove or refute. OB
    Structured/tabular ML in group theory he2023learning Supervised learning (classifier/regressor) Produces conjectures or predictions from engineered representations of algebraic objects; requires mathematical validation. OB
    Formal proof engineering (assistance within an ITP ecosystem)
    ImProver (Lean) ahuja2024improver Transformer (LLM program synthesis) Edits or refactors Lean proofs under constraints (e.g., length), helping with proof maintenance and presentation. IF
    • User understanding: BB = black-box use; IF = interface or formalism (e.g., Lean, JGEX); PL = pipeline configuration and objective (representation, search, scoring); OB = feature or representation choices and the validation procedure used to interpret outputs.
    Table 1: Highly specialized (“narrow”) systems cited in this paper, summarized by architecture class, primary mathematical role, and the typical user understanding required.
@@
    General purpose, “broad” models that can interact with humans via natural language, understand input images, and use tools (being able to call, in particular, specialized models, which can be viewed as a particular kind of “tool”, as we argue below), can provide general assistance to a mathematician (both with routine tasks, such as literature search, and domain-dependent tasks, such as counterexample search) and are not tied to a specific area of mathematics. These are what we refer to as “Mathematical Copilots”, as their goal is not simply to provide correct proofs but also to enhance the user’s understanding of the results and guide them towards their own discoveries (collins2023evaluating). Currently, the best candidates for these systems are large language models (LLMs), which we take to include multi-modal models that also admit images as components of the prompt. We contend that LLMs represent early predecessors of future general-purpose models (mathematical copilots), ultimately acting as mathematicians’ “thought partners” (collins2024building). Anecdotal evidence of such usefulness of current LLMs has been presented in various discussion forums, [^2] https://www.reddit.com/r/math/comments/14p6j5c/terence_tao_on_using_gpt4_to_help_with_math/ while noting the failures that LLMs often succumb to. However, there are concrete ways in which LLMs can be improved to become mathematical copilots: As indicated by the tool-integrated-reasoning approach, championed by models such as ToRA (gou2023tora) or Numina, [^3] https://huggingface.co/AI-MO/NuminaMath-7B-TIR general LLMs can incorporate other symbolic systems to which they can delegate symbolic tasks to. The mixture-of-experts paradigm (cai2024survey), on the other hand, highlights different ways in which LLMs can incorporate other LLMs as specially trained “modules” to which they can delegate mathematical tasks that are not symbolic in nature but still require particular mathematical skills. Such a hierarchical system can help mathematicians both on routine academic tasks (see (frieder2023large) for a short, high-level overview) and specialized ones. We envision that such models will require significantly less effort from users to use them, compared to specialized models, being self-explanatory, where natural language is the main mode of interaction. (While the outlined architectural approaches provide a viable step forward, in some cases, no data exists on which to train these models; see below.)
@@
    Fully automated, “universal” models that can generate mathematical theorems and theories autonomously, merely by being directed to a mathematical result. We envision that these systems will be the true successors of automated theorem provers (ATPs) such as the Vampire (riazanov2002design; kovacs2013first) or E (schulz2002brainiac) ATPs. These systems have not been widely adopted by the mathematical community. While there have been attempts to imbue these systems with (non-LLM) machine learning techniques (holden2021ML4ATP), these have not yet dramatically increased their performance. Nonetheless, once general-purpose models have been established, it is conceivable one could put these in automation loops to obtain “AI mathematicians”. For the domain of AI, this has already been recently investigated in the form of an “AI Scientist” (lu2024ai) that autonomously generated machine learning research articles. These are systems that conceivably will require minimal human input, merely guiding them towards certain theorems that one wants to see analyzed (proved, refuted, with comments on whether slightly different variations of a theorem, with slightly altered hypotheses, admit a proof).
We focus in this article on general purpose models, as outlined above, which are mathematical copilots. LLMs are currently the systems that hold the most promise to become, given a better data foundation, general purpose models. Fully automated, universal models are currently out of reach, as they are contingent on the former category of models. These are the systems that are ultimately sought when one speaks of “AI Mathematicians” (bengio2024machine). However, without the proper datasets, it is unlikely that these will come into existence. Hence, our focus is on the data that is needed to arrive at general-purpose models, the mathematical copilots. We have deliberately avoided drawing a hard boundary among the properties a model must fulfill to be placed in one of these three because the boundaries will be fluid, as the levels of automation and required human interaction are continuous scales.
@@
Several general principles have emerged from decades of machine learning research that shape expectations about AI progress. Perhaps the most influential is Sutton’s “bitter lesson” (sutton2019bitter), which observes that general methods leveraging computation (particularly search and learning) consistently outperform approaches that encode human domain knowledge. More recently, researchers have argued that open-endedness – the capacity for systems to continually generate novel, learnable artifacts – is essential for artificial superhuman intelligence (hughes2024openendedness). However, there are reasons to believe these principles may not straightforwardly transfer to mathematical AI. Mathematics demands logically rigorous proofs that must be verifiable, not merely plausible; pattern matching and statistical regularities, which suffice in domains like vision or game-playing, are insufficient for mathematical reasoning. Moreover, the “bitter lesson” presupposes abundant data that scales with computation, yet high-quality mathematical data, particularly at research level, is scarce, and much mathematical knowledge is implicit in workflows and heuristics that textbooks do not record. Open-endedness in mathematics requires more than generating novel outputs: it requires outputs that are correct and meaningful within a precise logical framework. These considerations suggest that progress in AI for mathematics may require domain-specific innovations in data, evaluation, and methodology that go beyond what has succeeded elsewhere in machine learning.
The limits of end-to-end generation for mathematics.
A recurring theme in machine learning is that end-to-end approaches – where models learn directly from inputs to outputs without explicit intermediate structure – often outperform systems with hand-designed pipelines (huang2023mlagentbench). However, we argue that for mathematical reasoning, pure end-to-end generation must be abandoned. Mathematics demands proofs that are not merely plausible but logically valid, and errors in intermediate steps propagate to invalidate entire arguments. Recent reasoning models address this by incorporating explicit chain-of-thought processes and, in formal settings, by interacting with proof assistants that verify each step (first2023baldur; song2024towards). The success of these hybrid approaches, which combine neural generation with symbolic verification, suggests that mathematical AI requires structured reasoning modes rather than monolithic sequence-to-sequence generation.
Our main reader audience consists of machine learning researchers, as we intend this article to raise awareness of these issues. Automated theorem provers (ATPs) and interactive theorem provers (ITPs), which have a decades-old history (harrison2014history), also had the ambition of realizing the goal of implementing mathematical copilots. Unfortunately, history has shown that if the computer science community and the mathematical community do not operate in lockstep, there is a risk of obtaining systems that will not be of interest to practicing research mathematicians; maybe controversially, blanchette2012more mentions that certain efforts to make these systems more user-friendly have perhaps reached a “plateau”. (Nonetheless, ATPs and ITPs have, in other domains, contributed significantly to driving the field of software automation forward.)
@@ 2 An Overview of Data-Related Issues
Without using tools specialized for mathematics, the current general models, LLMs, simply by well-designed training methodologies and trained on mathematical data, have demonstrated unprecedented capabilities in generating humanlike mathematical text, solving complex problems, and even engaging in creative problem-solving. For example, GPT-4 has performed promisingly on undergraduate-level mathematics formulated in natural language, as pointed out by frieder2023mathematical. Math-Specialized Gemini 1.5 Pro (reid2024gemini), a commercial model by Google not available to the public, has been reported to have an accuracy of over 90% on the MATH dataset (hendrycks2021measuring), one of the most widely-used datasets for testing mathematical ability. This score has recently been replicated by an open-weight model, QwQ. [^4] https://qwenlm.github.io/blog/qwq-32b-preview/ Attaining such a high score is equivalent to achieving the ability of an IMO gold medallist (according to (hendrycks2021measuring)). However, to date, no reports have been made where the mathematical benchmarks of QwQ are tested for contamination; not undertaking tests may lead to this score not being reproducible on a different dataset from the same difficulty and problem type distribution (yang2023rethinking; xu2024benchmarking).
Benchmark saturation on MATH.
Recent reasoning models approach or exceed 90% accuracy on the MATH benchmark (hendrycks2021measuring), which was considered extremely challenging at introduction (models scored below 7%). On MATH-500, a commonly used 500-problem subset, DeepSeek-R1 reports 97.3% deepseek2025r1 and QwQ-32B-Preview reports 90.6% qwen2024qwq32bpreview. Earlier models scored lower: Gemini 1.5 Pro achieved 77.9% on full MATH with majority voting reid2024gemini. This rapid progression leaves limited headroom under boxed-answer scoring and motivates harder benchmarks like FrontierMath (glazer2024frontiermath).
DeepTheorem reports a large corpus of informal IMO-style theorems paired with proofs, and uses entailed and contradictory variants to provide a binary reward signal for reinforcement learning in informal theorem proving (deeptheorem2025). Notably, this is a step beyond standard boxed-answer scoring: the supervision and evaluation are tied to logical consistency across theorem variants rather than only matching a numeric final answer.
These benchmark improvements do not, by themselves, resolve the evaluation and data problems that matter for mathematical copilots. First, high accuracy under boxed-answer scoring (still common in public benchmarks) leaves ambiguity about what changed, since multiple failure modes in the reasoning trace can be invisible once only the final answer is checked. Second, evidence from out-of-distribution tests suggests that strong benchmark scores can coexist with brittle reasoning on simple variants of problems (nezhurina2024alice). Third, as scores rise, contamination and memorization become harder to rule out with simple string-based decontamination, especially when mathematically equivalent problems can be phrased in many surface forms (xu2024benchmarking). These points motivate evaluation sets with stronger verification and a shift toward datasets that expose intermediate structure.
Further notable general-purpose LLMs are open-weight models with strong reported performance on baseline datasets, such as MATH and GSM8K (cobbe2021training), the DeepSeek family of models (shao2024deepseekmath; liu2024deepseek; zhu2024deepseek; xin2024deepseek), and the Qwen family of models (bai2023qwen; yang2024qwen2; hui2024qwen2; yang2024qwen2math). Other LLMs focus more on specific abilities such as MathPrompter (imani2023mathprompter), which associates a confidence value to arithmetic problem solutions, MathVista (lu2023mathvista), which solves geometric reasoning problems, WizardMath (luo2023wizardmath), or Llemma (azerbayev2023llemma), which was specialized through further pretraining on a math corpus, and is both open-weight and open-source. A special use case where LLMs have found success is in interactive theorem provers (ITPs) such as Isabelle (nipkow2002isabelle) or Lean (moura2021lean4), where they are used to prove theorems most prominently in the context of generating proofs of given formal statements (first2023baldur; zheng2023lyra; xin2023lego), autoformalization (szegedy2020promising; jiang2022draft), or providing code snippets of formal mathematics (song2024towards). For a survey of deep learning approaches on both formal and informal (natural language) mathematics, we refer to lu2022survey, and for more recent information on language models, including their performance on mainstream datasets, we refer to zhao2023survey.
@@
This paper argues that the current dataset landscape does not support the advancement of such general-purpose AI systems, in particular LLMs, towards a level of mathematical performance that makes them usable as daily mathematical thought partners (collins2024building) that help mathematicians push the boundaries of what is known, and capture the richness of mathematical reasoning and invention that human mathematical minds are capable of (zhang2023cogsciMath; dehaene2011number; feigenson2004core). Our focus thus includes, in particular, research-level mathematics, which involves many steps beyond deriving or formalizing proofs – the aspects of mathematical practice that are currently most strongly represented in data.
@@
The spectrum of tasks a copilot can assist with is not solely related to mathematics: programmers presently carry out programming tasks, such as fixing bugs or other GitHub issues (jimenez2024swebench), using various code-generating copilots, such as the GitHub Copilot (chen2021evaluating); the emerging Lean Copilot ecosystem (yang2023leandojo; song2024largelanguagemodelscopilots) assists users in the task of formalizing a mathematical theorem; Tutor CoPilot (wang2024tutor) provides real-time guidance to human tutors that teach school students mathematics at the level of math word problems.
Limits of relying on a single formal system.
Interactive theorem provers such as Lean provide machine-checkable correctness, and stronger automation inside these systems is valuable. However, there are limits to what is natural to express and maintain inside a single formal foundation and library, even when the underlying mathematics is standard. The following non-exhaustive examples illustrate why a mathematical copilot cannot be reduced to a Lean-only workflow.
First, formal developments can incur large overhead when a concept has many surface variants. Patrick Massot has pointed out this issue using the elementary calculus notion of a limit: a naive approach that treats each variant of “limit at a point” as a distinct definition leads to a combinatorial explosion in the number of definitions and routine plumbing lemmas (his talk gives concrete counts and examples) (massotLimitsTalk). In practice, mathematicians avoid learning and manipulating hundreds of separate limit definitions by teaching an abstract generator behind them, or by moving to a more uniform formalism (for example, ultrafilter-based formulations). Both approaches can be implemented formally, but they change what is convenient to teach, state, and reuse.
Second, some common mathematical activities are more naturally carried out in informal language. Concise proof exposition, proof summarization, and “zooming out” to communicate the core idea of a long argument are straightforward in natural language, while formal proofs typically expose low-level structure. Producing readable summaries from formal developments is a separate task that requires its own representations and data.
Third, conjecturing and exploration tend to happen before the final formal statement is stable. Early-stage work often involves proposing definitions, testing variants, and revising hypotheses. Doing all of this directly in Lean can add technical overhead at a point where the main goal is to move quickly through candidate formulations.
Fourth, the choice of foundations affects what counts as “the same” object. For example, work in homotopy type theory was driven in part by a refined analysis of equality and equivalence. Informal mathematics routinely treats objects as equal “at the level that matters” in a given theory, even when they are not definitional equalities in a fixed foundational encoding. Formal developments must commit to a specific notion of equality and transport, and that commitment can make some styles of reasoning less direct.
In the course of proving a theorem, one may need to convert between more abstract mathematics and less abstract mathematics. Flexible conversion might prove to be more difficult in Lean.
These limitations do not argue against formalization. They suggest that Lean is one tool in a broader toolbox, and that mathematical copilots will likely need a federation of tools and representations, including automation over informal mathematics, to support rapid iteration in research workflows.
Mathematics, with its rich set of domain-dependent workflows, tools, and modes of thinking, requires copilots with exceptionally broad capabilities. A mathematical copilot that is useful to a mathematician must significantly exceed the capabilities of the three exemplary copilots mentioned above, which are systems that are designed to help with a narrow technical set of tasks.
@@
Training corpora vs benchmarks.
In this paper, the term dataset covers two use cases: (i) corpora used for pretraining, fine-tuning, or reinforcement learning, and (ii) held-out benchmarks used to report model performance. Some collections are used in both ways. Benchmarks are also sometimes converted into training data, for example by creating instruction-style variants of benchmark problems. As a result, many of the issues discussed below apply to both training and benchmarking, but they affect different parts of the pipeline. Difficulty calibration and interaction-mode coverage affect the supervision signal during training and the amount of headroom in a benchmark. Contamination affects both use cases: benchmark items can appear in training corpora, and benchmark scores can be inflated when test problems, or close variants, leak into training data. Other issues depend more strongly on the use case. Binary scoring and the evaluation protocol determine what a benchmark score measures. For formal mathematics, placeholders (sorry, Admitted), duplication across alternative formalizations, and toolchain provenance affect both the quality of training data and the interpretation of evaluation results. Table 2 summarizes these distinctions.
To give an overview of our contributions in that section, current natural language datasets typically model a question-response interaction, such as posing a problem and writing a solution. We identify dimensions where such datasets frequently fall short of being able to assess advanced mathematics:
@@
  • Standardized Interaction Modes: Often, natural language datasets only contain question-answer mathematical problems. The daily experience of a mathematician is much richer than that (zhang2023cogsciMath), since to prove/search for a theorem, all kinds of “lateral” modes of thinking have to be applied. Examples are intermediate conjecture formulation, (counter)example search, and finding analogies with other theorems. These are typically not found in textbooks, so there is no representation of these in terms of data. We point to collins2023evaluating for an exposition of the problems of static evaluation for natural-language mathematics problems in practice. In short, static evaluation relies on fixed input–output pairs that fail to capture iterative or adaptive interactions, such as clarifying questions and corrections, that are an integral part of effective problem solving.
@@
  • (Binned) Sankey Diagrams: Most benchmarks are snapshots of current models, and don’t record clearly how performance is changing. This makes it hard to assess the broader direction of evolution of language models w.r.t. mathematics. (Binned) Sankey Diagram offer an easy remedy, but have not been embraced by the community.